### v3

# Your Student is Better Than Expected: Adaptive Teacher-Student Collaboration for Text-Conditional Diffusion Models

Abstract
Knowledge distillation methods have recently shown to be a promising direction to speedup the synthesis of large-scale diffusion models by requiring only a few inference steps. While several powerful distillation methods were recently proposed, the overall quality of student samples is typically lower compared to the teacher ones, which hinders their practical usage. In this work, we investigate the relative quality of samples produced by the teacher text-to-image diffusion model and its distilled student version. As our main empirical finding, we discover that a noticeable portion of student samples exhibit superior fidelity compared to the teacher ones, despite the “approximate” nature of the student. Based on this finding, we propose an adaptive collaboration between student and teacher diffusion models for effective text-to-image synthesis. Specifically, the distilled model produces an initial image sample, and then an oracle decides whether it needs further improvements with the teacher model. Extensive experiments demonstrate that the designed pipeline surpasses state-of-the-art text-to-image alternatives for various inference budgets in terms of human preference. Furthermore, the proposed approach can be naturally used in popular applications such as text-guided image editing and controllable generation.
Figure 1: Left: Overview of the proposed approach. Right: Side-by-side comparison of SDv1.51.51.51.5 and SDXL with their few-step distilled versions. The distilled models surpass the original ones in a noticeable number of samples for the same text prompts and initial noise.
1 Introduction
Large-scale diffusion probabilistic models (DPMs) have recently shown remarkable success in text-conditional image generation [34, 38, 41, 32] that aims to produce high quality images closely aligned with the user-specified text prompts. However, DPMs pose sequential synthesis leading to high inference costs opposed to feed-forward alternatives, e.g., GANs, that provide decent text-to-image generation results for a single forward pass [43, 17].
There are two major research directions mitigating the sequential inference problem of state-of-the-art diffusion models. One of them considers the inference process as a solution of a probability flow ODE and designs efficient and accurate solvers [49, 18, 26, 27, 58] reducing the number of inference steps down to ${\sim}10$ without drastic loss in image quality. Another direction represents a family of knowledge distillation approaches [42, 50, 30, 24, 25, 28, 12, 44] that learn the student model to simulate the teacher distribution requiring only $1{-}4$ inference steps. Recently, distilled text-to-image models have made a significant step forward [30, 25, 28, 44]. However, they still struggle to achieve the teacher performance either in terms of image fidelity and textual alignment [30, 25, 28] or distribution diversity [44]. Nevertheless, we hypothesize that text-to-image students may already have qualitative merits over their teachers. If so, perhaps it would be more beneficial to consider a teacher-student collaboration rather than focusing on replacing the teacher model entirely.
In this paper, we take a closer look at images produced by distilled text-conditional diffusion models and observe that the student can generate some samples even better than the teacher. Surprisingly, the number of such samples is significant and sometimes reaches up to half of the empirical distribution. Based on this observation, we design an adaptive collaborative pipeline that leverages the superiority of student samples and outperforms both individual models alone for various inference budgets. Specifically, the student model first generates an initial image sample given a text prompt, and then an “oracle” decides if this sample should be updated using the teacher model at extra compute. The similar idea has recently demonstrated its effectiveness for large language models (LLMs) [5] and we show that it can be naturally applied to text-conditional diffusion models as well. Our approach is schematically presented in Figure 1. To summarize, our paper presents the following contributions:
  • We reveal that the distilled student DPMs can outperform the corresponding teacher DPMs for a noticeable number of generated samples. We demonstrate that most of the superior samples correspond to the cases when the student model significantly diverges from the teacher.
  • Based on the finding above, we develop an adaptive teacher-student collaborative approach for effective text-to-image synthesis. The method not only reduces the average inference costs but also improves the generative quality by exploiting the superior student samples.
  • We provide an extensive human preference study illustrating the advantages of our approach for text-to-image generation. Moreover, we demonstrate that our pipeline can readily improve the performance of popular text-guided image editing and controllable generation tasks.
2 Related work
Diffusion Probabilistic Models (DPMs) [14, 48, 49] represent a class of generative models consisting of forward and reverse processes. The forward process $\{\boldsymbol{x}_{t}\}_{[0,T]}$ transforms real data $\boldsymbol{x}_{0}\sim p_{\text{data}}(\boldsymbol{x}_{0})$ into the noisy samples $\boldsymbol{x}_{t}$ using the transition kernels $\mathcal{N}\left(\boldsymbol{x}_{t}\ |\ \sqrt{1-\sigma_{t}}\boldsymbol{x}_{0},\sigma_{t}\mathbf{I}\right)$ specifying $\sigma_{t}$ according to the selected noise schedule.
The reverse diffusion process generates new data points by gradually denoising samples from a simple (usually standard normal) distribution. This process can be formulated as a probabilistic-flow ODE (PF-ODE) [46, 49], where the only unknown component is a score function, which is approximated with a neural network. The ODE perspective of the reverse process fosters designing a wide range of the specialized solvers [26, 58, 27, 46, 18, 57, 15] for efficient and accurate sampling. However, for text-to-image generation, one still needs ${\sim}25$ and more steps for the top performance.
Text-conditional diffusion models can be largely grouped into cascaded and latent diffusion models. The cascaded models [41, 32] generate a sample in several stages using separate diffusion models for different image resolutions. The latent diffusion models [34, 38, 37] first generate a low-resolution latent variable in the VAE [21] space and then apply its feedforward decoder to map the latent sample to the high-resolution pixel space. Thus, the latent diffusion models have significantly more efficient inference thanks to a single forward pass for the upscaling step.
Figure 2: Student outperforms its teacher (SD1.5). Left: Text-conditional image synthesis. Right: Text-guided image editing (SDEdit [29]). The images within each pair are generated for the same initial noise sample.
Along with cascaded models, there are other works combining several diffusion models into a single pipeline. Some methods propose to use distinct diffusion models at different time steps [1, 23, 8, 55]. Others [34, 25] consider an additional model to refine the samples produced with a base model. In contrast, our method relies on the connection between student and teacher models and adaptively improves only selected student samples to reduce the inference costs.
Text-to-image diffusion models have also succeeded in text-guided image editing and personalization [3, 29, 31, 19, 40, 10]. Moreover, some methods allow controllable generation via conditioning on additional inputs, e.g., canny-edges, semantic masks, sketches [56, 52]. Our experiments show that the proposed pipeline is well-suited to these techniques.
Distillation of diffusion models is another pivotal direction for efficient diffusion inference [50, 47, 42, 30, 2, 25, 44]. The primary goal is to adapt the diffusion model parameters to represent the teacher image distribution for $1{-}4$ steps. Recently, consistency distillation (CD) [50] have demonstrated promising results on both classical benchmarks [20, 47] and text-to-image generation [28] but fall short of the teacher performance at the moment. Concurrently, adversarial diffusion distillation [44] could outperform the SDXL-Base [34] teacher for $4$ steps in terms of image quality and prompt alignment. However, it significantly reduces the diversity of generated samples, likely due to the adversarial training [11] and mode-seeking distillation technique [35]. Therefore, it is still an open question if a few-step distilled model can perfectly approximate the diffusion model on highly challenging and diverse distributions that are currently standard for text-conditional generation [45].
3 Toward a unified teacher-student framework
Opposed to the purpose of replacing the expensive text-to-image diffusion models by more effective few-step alternatives, the present work suggests considering the distilled text-to-image models as a firm companion in a teacher-student collaboration.
In this section, we first explore the advantages of the distilled text-to-image models and then unleash their potential in a highly effective generative pipeline comprising the student and teacher models.
3.1 Delving deeper into the student performance
We start with a side-by-side comparison of the student and teacher text-to-image diffusion models. Here, we focus on Stable Diffusion v1.5111https://huggingface.co/runwayml/stable-diffusion-v1-5 (SD1.51.51.51.5) as our main teacher model and distill it using consistency distillation [50]. The student details and sampling setup are presented in A. The similar analysis for a few other distilled models is provided in B.2.
Figure 3: (a) Visual examples of similar (Left) and dissimilar (Right) teacher and student samples. (b) Similarity between the student and teacher samples w.r.t. the difference in sample quality. Highly distinct samples tend to be of different quality. (c) Human vote distribution for different distance ranges between student and teacher samples. Most of the student wins are achieved when the student diverges from the teacher.
Figure 4: Effect of image complexity. (a) More similar student and teacher samples corresponds to simpler images and vice versa. (b) The student and teacher largely diverge in image quality on the complex teacher samples.
Figure 5: Effect of text prompts. (a) Shorter prompts usually lead to more similar student and teacher samples. (b) The student and teacher tend to generate more similar images when the student relies heavily on the text prompt.
In Figure 1 (Right), we provide the human votes for 600 random text prompts from COCO2014 [22] for SD1.51.51.51.5 and SDXL. The images within each pair are generated for the same initial noise sample. We observe that the students generally falls short of the teacher performance. However, interestingly, despite the initial intention to mimic the teacher model, ${\sim}30\%$ student samples were preferred over the teacher ones. A few visual examples in  Figure 2 validate these results. Therefore, we can formulate our first observation:
The student can surpass its teacher in a substantial portion of image samples.
Below, we develop a more profound insight into this phenomenon.
Student-teacher similarity. First, we evaluate the student’s ability to imitate the teacher. We compute pairwise distances between the student (S) and teacher (T) images generated for the same text prompts and initial noise. As a distance measure, we use DreamSim [9] tuned to be aligned with the human perception judgments. For evaluation, we consider 5000 prompts from the COCO2014 [22] validation split.
Primarily, we observe that many student samples are highly distinct from the teacher ones. A few image pairs are presented in Figure 3a. Figure 3c presents the human vote distribution for low ($0{-}20\%$), medium ($40{-}60\%$) and high ($80{-}100\%$) distance ranges. Interestingly, most of the student wins appear when its samples are highly different from the teacher ones. This brings us to our second observation:
The student wins are more likely where its samples significantly differ from the teacher ones.
Also, we evaluate the relative gap in sample quality against the similarity between the teacher and student outputs. To measure the quality of individual samples, we use ImageReward [53], which shows a positive correlation with human preferences in terms of image fidelity and prompt alignment. The divergence in quality is calculated as the difference between the ImageReward scores for student and teacher samples. We observe that highly distinct samples likely have a significant difference in quality. Importantly, this holds for both student failures and successes, as shown in Figure 3b. Therefore, effectively detecting the positive student samples and improving the negative ones can potentially increase the generative performance.
Image complexity. Then, we describe the connection of the similarity between student and teacher samples with the teacher image complexity. To estimate the latter, we use the ICNet model [7] learned on a large-scale human annotated dataset. The results are presented in Figure 4. We notice that larger distances between student and teacher outputs are more typical for complex teacher samples. In other words, the student mimics its teacher for plain images, e.g., close-up faces, while acting more as an independent model for more intricate ones. Figure 4b confirms that significant changes in image quality are observed for more complex images.
Text prompts. Then, we analyse the connection of the student-teacher similarity with the prompt length. Figure 5 demonstrates that shorter prompts typically lead to more similar teacher and student samples. Here, the prompt length equals to the number of CLIP tokens. Intuitively, longer prompts are more likely to describe intricate scenes and object compositions than shorter ones. Note that long prompts can also carry low textual informativeness and describe concepts of low complexity. We hypothesize that this causes high variance in Figure 5a.
Also, we report the prompt influence on the student generation w.r.t. the student-teacher similarity in Figure 5b. We estimate the prompt influence by aggregating student cross-attention maps. More details are in B.1. The student tends to imitate the teacher if it relies heavily on the text prompt.
Trajectory curvature. Previously, it was shown to be beneficial to straighten the PF-ODE trajectory before distillation [24, 25]. We investigate the effect of the trajectory curvature on the similarity between the teacher and student samples and its correlation with the teacher sample complexity. We estimate the trajectory curvatures following [4] and observe that straighter trajectories lead to more similar student and teacher samples (Figure 6a). In addition, we show that the trajectory curvature correlates with the teacher sample complexity (Figure 6b).
Figure 6: Effect of teacher trajectory curvature. (a) The student samples resemble the teacher ones for less curved trajectories. (b) Straighter trajectories usually correspond to plainer teacher images.
Figure 7: Full-reference vs no-reference decision-making. Usually, one can find a $k$-th percentile of the ImageReward scores providing the correlation with human votes similar to the full-reference comparisons but without observing the teacher samples.
To sum up, we conclude that the student largely diverges from the teacher on the samples that are challenging in different respects. Interestingly, the superior student samples often occur in these cases.
3.2 Method
In this section, we propose an adaptive collaborative approach consisting of three steps: 1) Generate a sample with the student model; 2) Decide if the sample needs further improvement; 3) If so, refine or regenerate the sample with the teacher model.
Student generation step produces an initial sample $\mathcal{X}^{\mathbf{S}}$ for a given context and noise. This work considers consistency distillation [50] as a primary distillation framework and uses multistep consistency sampling [50] for generation.
Adaptive step leverages our finding that many student samples may exhibit superior quality. Specifically, we seek an “oracle” that correctly detects superior student samples. For this role, we consider an individual sample quality estimator $\mathbf{E}$. In particular, we use the current state-of-the-art automated estimator, ImageReward (IR) [53] that is learned to imitate human preferences for text-to-image generation.
Then, comparing the scores of the teacher and student samples, one can conclude which one is better. However, in practice, we avoid expensive teacher inference to preserve the efficiency of our approach. Therefore, a decision must be made having access only to the student sample $\mathcal{X}^{\mathbf{S}}$. To address this problem, we introduce a cut-off threshold $\tau$ which is a $k$-th percentile of the IR score tuned on a hold-out subset of student samples. The details on the threshold tuning are described in C. During inference, the IR score is calculated only for $\mathcal{X}^{\mathbf{S}}$. If it exceeds the threshold $\tau$, we accept the sample and avoid further teacher involvement. Interestingly, we observe that it is often possible to reproduce the accuracy of the full-reference estimation by varying $\tau$ (see Figure 7). Also, note that IR calculation costs are negligible compared to a single diffusion step, see D.7.
Improvement step engages the teacher to improve the quality of the rejected student samples. We consider two teacher involvement strategies: regeneration and refinement. The former simply applies the teacher model to produce a new sample from scratch for the same text prompt and noise. The refinement is inspired by the recent work [34]. Specifically, $\mathcal{X}^{\mathbf{S}}$ is first corrupted with a Gaussian noise controlled by the rollback value $\sigma\in[0,1]$. Higher $\sigma$ leads to more pronounced changes. We vary $\sigma$ between $0.3$ and $0.75$ in our experiments. Then, the teacher starts sampling from the corrupted sample following the original noise schedule and using an arbitrary solver, e.g., DPM-Solver [26]. Note that refinement requires significantly fewer steps to produce the final sample than generation from scratch. Intuitively, the refinement strategy aims to fix the defects of the student sample. At the same time, the regeneration strategy may be useful if $\mathcal{X}^{\mathbf{S}}$ is poorly aligned with the text prompt in general. Our experiments below confirm this intuition.
Figure 8: Qualitative comparison of our adaptive refinement approach to the SD1.5 teacher model.
Figure 9: User preference study (SD1.5). (a) Comparison of our approach to the top-performing teacher configurations. (b) Comparison to the teacher model with DPM-Solver for the same average number of steps. (c) Comparison to the refinement strategy without the adaptive step for the same average number of steps. Top row: LAION-Aesthetic text prompts. Bottom row: COCO2014 text prompts. For our adaptive approach, we use the refinement strategy (R).
4 Experiments
We evaluate our approach for text-to-image synthesis, text-guided image editing and controllable generation. The results confirm that the proposed adaptive approach can outperform the baselines for various inference budgets.
4.1 Text-guided image synthesis
In most experiments, we use Stable Diffusion v1.5 (SD1.5) as a teacher model and set the classifier-free guidance scale to $8$. To obtain a student model, we implement consistency distillation (CD) for latent diffusion models and distill SD1.5 on the 80M subset of LAION2B [45]. The resulting model demonstrates decent performance for $5$ steps of multistep consistency sampling with the guidance scale $8$.
Metrics. We first consider FID [13], CLIP score [36] and ImageReward [53] as automatic metrics. ImageReward is selected due to a higher correlation with human preferences compared to FID and CLIP scores. OpenCLIP ViT-bigG [6] is used for CLIP score calculation. For evaluation, we use $5000$ text prompts from the COCO2014 validation set [22].
Also, we evaluate user preferences using side-by-side comparisons conducted by professional assessors. We select $600$ random text prompts from the COCO2014 validation set and $600$ from LAION-Aesthetics. More details on the human evaluation pipeline are provided in D.1.
Configuration. For our adaptive approach, we consider both refinement (R) and regeneration (G) strategies using a second order multistep DPM solver [27] and vary the number of sampling steps depending on the average inference budget. As a sample estimator $\mathbf{E}$, we consider ImageReward, except for the CLIP score evaluation. For each inference budget, we tune the hyperparameters $\sigma$ and $\tau$ on the hold-out prompt set. The exact values are provided in D.2.
Baselines. We consider the teacher performance as our main baseline and use DDIM [46] for 50 steps and a second order multistep DPM solver [27] for lower steps. In addition, we compare to the refining strategy on top of all student samples, without the adaptive step. This baseline is inspired by the recent results [34] demonstrating the advantages of the refinement stage itself. Also, we provide the comparison with Restart Sampling [54].
Figure 10: Automated evaluation (SD1.5). Comparison of the FID, CLIP and ImageReward scores for different number of sampling steps on 5K text prompts from COCO2014. The proposed collaborative approach outperforms all the baselines. The adaptive pipeline with the regeneration strategy (G) demonstrates higher textual alignment (CLIP score), while the refinement strategy (R) improves the image fidelity (FID).
Figure 11: User preference study (SDXL). Left: Comparison of the adaptive approach with CD-SDXL to the top-performing teacher setup. Right: Comparison of the adaptive approach with ADD-XL to ADD-XL for the same average number of steps.
Results. The quantitative and qualitative results are presented in Figures 9, 10 and Figure 12, respectively. According to the automatic metrics, our approach outperforms all the baselines. Specifically, in terms of CLIP scores, the adaptive regeneration strategy demonstrates superior performance compared to the refining-based counterpart. On the other hand, the adaptive refining strategy is preferable in terms of FID scores. We assume that the refinement strategy essentially improves the image fidelity and does not significantly alter the textual alignment due to the relatively small rollback values. In terms of ImageReward, both adaptive strategies perform equally.
[…]
SDXL results. In addition to SD1.5 experiments, we evaluate our pipeline using the recent CD-SDXL [28] and ADD-XL [44] which are both distilled from the SDXL-Base model [34]. Our approach with CD-SDXL stands against the top-performing SDXL setting: $50$ steps of the DDIM sampler. For ADD-XL, we provide the comparison for $4$ steps where ADD-XL has demonstrated exceptionally strong generative performance in terms of human preference [44]. In both settings, our approach uses the adaptive refinement strategy with the UniPC solver [58]. Note that both the SDXL-Base and SDXL-Refiner [34] models can be used for refinement. In our experiments, we observe that the refiner suits slightly better for fixing minor defects while the teacher allows more pronounced changes. Thus, we use the refiner for low $\sigma$ values and the base model for the higher ones. More setup details are provided in D.3.
[…]
In D.5, we also investigate how our approach affects the distribution diversity. D.4 aims to reveal the potential gains of our approach if the oracle accuracy increases in the future.
Method Steps DINOv2 $\downarrow$ ImageReward $\uparrow$
[…]
Table 1: Comparison of SDEdit using different approaches in terms of reference preservation and editing quality for the strength $0.6$.
Figure 12: Visual examples produced with our approach and the top-performing teacher (SD1.5) configuration. Top: Text-guided image editing with SDEdit [29]. Bottom: Controllable image generation with Canny edges and semantic segmentation masks using ControlNet [56].
4.2 Text-guided image editing
This section applies our approach for text-guided image editing using SDEdit [29]. We add noise to an image, alter the text prompt and denoise it using the student model first. If the edited image does not exceed the threshold $\tau$, the teacher model is used for editing instead. In the editing setting, we observe that the refinement strategy significantly reduces similarity with the reference image due to the additional noising step. Thus, we apply the regeneration strategy only.
In these experiments, the SD1.5 and CD-SD1.5 models are considered. As performance measures, we use ImageReward for editing quality and DINOv2 [33] for reference preservation. For evaluation, $100$ text prompts from COCO2014 are manually prepared for the editing task.
Results. Table 1 provides evaluation results for a SDEdit noising strength value $0.6$. The proposed method demonstrates a higher ImageReward score compared to the baselines with similar reference preservation scores. In addition, we present the performance for different editing strength values in Figure 13. Our approach demonstrates a better trade-off between reference preservation and editing quality. We provide qualitative results in Figure 12.
Figure 13: SDEdit performance for different strength values in terms of reference preservation (DINOv2) and editing quality (IR).
4.3 Controllable image generation
Finally, we consider text-to-image generation using Canny edges and semantic segmentation masks as an additional context and use ControlNet [56] for this task. We use ControlNet pretrained on top of SD1.5 and directly plug it into the distilled model (CD-SD1.5). Interestingly, the model pretrained for the teacher model fits the student model surprisingly well without any further adaptation.
For the teacher model, the default ControlNet sampling configuration is used: $20$ sampling steps of the UniPC [58] solver. In our adaptive approach, we use the refinement strategy with $10$ steps of the same solver. For performance evaluation, we conduct the human preference study for each task on $600$ examples and provide more details in D.6.
Results. According to the human evaluation, our approach outperforms the teacher ($20$ steps) by $19\%$ ($9$ steps) and $4\%$ ($11$ steps) for Canny edges and semantic segmentation masks, respectively. The visual examples are in Figure 12.
5 Conclusion
This work investigates the performance of the distilled text-to-image models and demonstrates that they may consistently outperform the teachers on many samples. We design an adaptive text-to-image generation pipeline that takes advantage of successful student samples and, in combination with the teacher model, outperforms other alternatives for low and high inference budgets.
Supplementary Material
Appendix A CD-SD1.5 implementation details
In this work, we develop consistency distillation (CD) for Stable Diffusion following the official implementation [50]. For training, we prepare a subset of LAION2B [45], which consists of $80$M image-text pairs. As a teacher sampler, we consider DDIM-solver using $50$ sampling steps and Variance-Preserving scheme [49]. We use the teacher UNet architecture as a student model and initialize it with the teacher parameters. Classifier-free guidance is applied to the distilled model directly without merging it into the model as done in [30]. During training, we uniformly sample the guidance strength from $1$ to $8$. Thus, our model supports different guidance scales during sampling. We train the student for ${\sim}200$K iterations on $8$ A100 GPUs using the following setting: $512$ batch size; $0.95$ EMA rate; $1\text{e}{-}5$ fixed learning rate; L2 uniformly weighted distillation loss calculated in the latent space of the VAE encoder. During inference, the multistep stochastic sampler [50] is used to generate images. In most of our experiments, we use $5$ sampling steps.
Note that we use our implementation of consistency distillation for SD because, when most experiments were conducted, there were no publicly available implementations.
Figure 14: SD1.5 samples of different complexity according to the ICNet model [7].
Figure 15: Left: Two examples of diffusion model trajectories with high and low curvatures. Right: Trajectory deviations according to [4].
Appendix B Analysis
B.1 Details
Image complexity. To calculate the image complexity, we use the recent ICNet model [7]. This model is learned on a large-scale human annotated dataset. Each image corresponds to a complexity score ranging from 0 (the simplest) to 1 (the most complex). In Figure 14, we provide examples of Stable Diffusion samples with the lowest and highest image complexity. More complex images usually depict multiple entities, often including people, and intricate backgrounds.
Text influence. We calculate the influence of a text prompt on student generation by using cross-attention between token embeddings and intermediate image representations. Following [51], we collect cross-attention maps for all diffusion steps and UNet [39] layers. Then, the average attention score is calculated for each text token. Finally, the highest value among all tokens is returned.
Trajectory curvature is estimated according to the recent work [4]. First, we calculate the trajectory deviations as L2 distance from the denoised prediction at a time step $t$ to the straight line passing through the denoising trajectory endpoints. The trajectory curvature corresponds to the highest deviation over all time steps.
In Figure 15 (Left), we visualize two denoising trajectories corresponding to high and low curvatures. We apply PCA [16] to reduce the dimensionality of the denoised predictions. In addition, Figure 15 (Right) demonstrates trajectory deviations for different time steps. The highest deviations typically occur closer to the end of the denoising trajectory.
B.2 Other distilled text-to-image models
Here, we conduct a similar analysis as in Section 3 for consistency distillation on Dreamshaper v7777 [28] and architecture-based distillation [^2] https://huggingface.co/docs/diffusers/using-diffusers/distilled_sd.
In contrast to the few-step distillation approaches [30, 50, 20, 24, 28], the architecture-based distillation removes some UNet layers from the student model for more efficient inference and trains it to imitate the teacher using the same deterministic solver and number of sampling steps.
In Figures 1920, we show that both methods can produce samples that significantly differ from the teacher ones for the same text prompt and initial noise. Then, Figures 2122 confirm that other observations in Section 3 remain valid for the Dreamshaper and architecture-based students as well.
Input: $\mathbf{E},\mathbf{S},\mathbf{T}-$ estimator, student, teacher;
$\mathcal{I}-$ input; $\sigma,\tau-$ rollback value and cut-off threshold.
$\hat{\mathcal{O}}=\mathbf{S}(\mathcal{I})$
  // Student prediction
1 if $\mathbf{E}(\hat{\mathcal{O}})<\tau$  then
       Strategy 1:
        // Refinement
2           $\hat{\mathcal{O}}_{\sigma}=\sqrt{1-\sigma}\cdot\hat{\mathcal{O}}+\sqrt{\sigma}\cdot\mathcal{Z},\ \mathcal{Z}\sim\mathcal{N}(0,1)$
3           $\hat{\mathcal{O}}=\mathbf{T}(\mathcal{I},\hat{\mathcal{O}}_{\sigma})$
       Strategy 2:
        // Regeneration
4           $\hat{\mathcal{O}}=\mathbf{T}(\mathcal{I})$
return $\hat{\mathcal{O}}$
Algorithm 1 teacher-student adaptive collaboration
Appendix C Cut-off threshold tuning
The cut-off threshold $\tau$ is used for the adaptive selection of student samples. It corresponds to the $k$-th percentile of the metric values calculated on validation student samples. $600$ and $300$ samples are generated for tuning on the COCO2014 and LAION-Aesthetics datasets, respectively. Note that the prompts used for tuning do not overlap with the test ones. Then, we calculate an individual score, e.g., IR score, for each validation student sample and select the percentile based on an average inference budget or target metric value. For example, suppose we select the percentile given a $15$ step budget and intend to perform $5$ student steps and $20$ steps for the improvement strategy. In this case, we have to select $\tau$ as a $50$-th percentile, which results in the final average number of steps: $5+0.5\cdot 20=15$.
During inference, we perform the adaptive selection as follows: if the score of the student sample exceeds $\tau$, we consider that this sample might be superior to the teacher one and keep it untouched. Otherwise, we perform an improvement step using the teacher model (refinement or regeneration). The proposed pipeline is presented in Algorithm 1.
We also show that hyperparameter tuning is straightforward and requires negligible effort. To verify this, we tune the threshold using the various number of prompts (Tab. 2). We can see that $500$ prompts are sufficient for the threshold convergence. Thus, the process only needs to generate $500$ student samples and takes ${\sim}2.5$ minutes.
Prompts $150$ $250$ $500$ $5000$ $25000$
[…]
Table 2: The threshold value tuned using prompts from Diffusion-DB and PickScore datasets and the time required for tuning.
Appendix D Experiments
D.1 Human evaluation
To evaluate the text-to-image performance, we use the side-by-side comparison conducted by professional annotators. Before the evaluation, all annotators pass the training and undergo the preliminary testing. Their decisions are based on the three factors: textual alignment, image quality and aesthetics (listed in the priority order). Each side-by-side comparison is repeated three times by different annotators. The final result corresponds to the majority vote.
D.2 Experimental setup (SD1.5)
The exact hyperparameter values and number of steps used for the automated estimation (FID, CLIP score and ImageReward) of the adaptive refinement strategy in Table 3. The adaptive regeneration uses the same cut-off thresholds and number of steps, but the rollback value, $\sigma$, is equal to $1$. The values used for human evaluation are presented in Table 4.
D.3 Experimental setup (SDXL)
We evaluate two distilled models: CD-SDXL [28] and ADD-XL [44]. For the first evaluation, we use $4$ steps of the CD-SDXL and then apply $12$ adaptive refinement steps using the teacher model (SDXL-Base [34]) with the UniPC solver [58]. We compare our pipeline to the default teacher configuration: $50$ DDIM steps. For the second evaluation, we perform $2$ steps of the ADD-XL and $4$ steps of the SDXL-Refiner [34] for the adaptive refinement strategy. We compare to $4$ ADD-XL steps as this setting outperformed SDXL-Base in terms of image quality and textual alignment [34]. The exact hyperparameter values and number of steps used for human evaluation are in Table 5.
We found that SDXL-Refiner performs slightly better than the base model for small refinement budgets (e.g., $4$). The refiner typically helps to improve fine-grained details, e.g., face attributes or background details. However, it faces difficulties in providing global changes and sometimes brings artifacts for large rollback values, $\sigma$. Thus, we use the SDXL-Base teacher for more refinement steps (e.g., 12).
Metric $\sigma$ $k$ Steps
CD Refinement Adaptive
ImageReward $0.4$ $60$ $5$ $5$ $8$
$0.55$ $60$ $5$ $10$ $11$
$0.7$ $60$ $5$ $15$ $14$
$0.7$ $60$ $5$ $25$ $20$
$0.7$ $60$ $5$ $35$ $26$
$0.7$ $60$ $5$ $45$ $32$
CLIP score $0.5$ $60$ $5$ $5$ $8$
$0.7$ $60$ $5$ $10$ $11$
$0.75$ $60$ $5$ $15$ $14$
$0.75$ $60$ $5$ $25$ $20$
$0.75$ $60$ $5$ $35$ $26$
$0.75$ $60$ $5$ $45$ $32$
FID $0.75$ $40$ $3$ $5$ $5$
$0.7$ $70$ $3$ $15$ $14$
Table 3: Hyperparameter values used for the automated evaluation (SD 1.5), Figure 10.
Metric $\sigma$ $k$ Steps
CD Refinement Adaptive
Human evaluation $0.7$ $50$ $5$ $10$ $10$
$0.7$ $50$ $5$ $20$ $15$
$0.7$ $50$ $5$ $40$ $25$
Table 4: Hyperparameter values used for the user preference study (SD 1.5), Figure 9.
Metric $\sigma$ $k$ Steps
ADD-XL Refinement Adaptive
Human evaluation $0.4$ $50$ $2$ $4$ $4$
CD-SDXL Refinement Adaptive
Human evaluation $0.85$ $70$ $4$ $12$ $13$
Table 5: Hyperparameter values used for the user preference study (SDXL), Figure 11.
D.4 Effect of oracle accuracy
Figure 16: User preferences for different accuracy levels of the no-reference decision-making procedure. the automated sample estimator. Current state represents the results using ImageReward. The results for higher accuracy rates demonstrate the future gains if the oracle performance improves.
The potential bottleneck of our approach is a poor correlation of existing text-to-image automated estimators with human preferences. For example, ImageReward usually exhibits up to $65\%$ agreement with annotators. Moreover, it remains unclear what oracle accuracy can be achieved with no-reference decision-making, even if the estimator provides the perfect agreement. In Figure 16, we conduct a synthetic experiment examining the effect of the oracle accuracy on our scheme performance to reveal its future potential. We compare the adaptive refinement method ($10$ steps) to SD1.5 ($50$ steps) manually varying the oracle accuracy. We observe significant future gains even for the $75\%$ accuracy rate.
D.5 Distribution diversity
Figure 17: Diversity human scores collected for different methods.
Figure 18: Image complexity of the CD-SD1.5 and SD1.5 samples in terms of ICNet [7]. Left: Box plot representing the complexity quantiles for both models. Right: Distribution of individual complexity values. Each dot corresponds to a pair of samples generated for the same prompt and initial noise. The distilled model only slightly simplifies the teacher distribution.
In the proposed teacher-student collaboration, the oracle aims to accept high-quality and well-aligned student samples but does not control the diversity of the resulting image distribution. Therefore, if the student exhibits severe mode collapse or oversimplifies the teacher samples, the adaptive pipeline will likely inherit these issues to some extent.
In this section, we investigate this potential problem for several existing distilled text-to-image models. Specifically, we consider consistency distillation [28] for SD1.5 and SDXL [34] models and ADD-XL [44]. Note that ADD-XL is a GAN-based distillation method that generates exceptionally realistic samples but has evidence to provide poor image diversity for the given text prompt [44].
We estimate the diversity of generated images by conducting a human study. In more detail, given a text prompt and a pair of samples generated from different initial noise samples, assessors are instructed to evaluate the diversity of the following attributes: angle of view, background, main object and style. For each model, the votes are collected for $600$ text prompts from COCO2014 and aggregated into the scores from 0 to 1, higher scores indicate more diverse images. The results are presented in Figure 17.
CD-SDXL demonstrates significantly better diversity than ADD-XL but still produces less various images compared to the SDXL teacher. CD-SD1.5 performs similarly to the SD1.5 teacher. Also, both adaptive strategies increase the diversity of the SDXL student models, especially the regeneration one. In Figure 30, we illustrate the diversity of images generated with different distilled text-to-image models (ADD-XL, CD-SDXL and CD-SD1.5). Each column corresponds to a different initial noise (seed). We notice that ADD-XL exhibits the lowest diversity compared to the CD-based counterparts.
Then, we address whether the distilled models tend to oversimplify the teacher distribution. In this experiment, we evaluate SD1.5 using DDIM for $50$ steps and the corresponding CD-SD1.5 using $5$ sampling steps. In Figure 18, we compare the complexity of the student and teacher samples in terms of the ICNet score [7]. We observe that CD-SD1.5 imperceptibly simplifies the teacher distribution.
To sum up, in our experiments, the CD-based models provide the decent distribution diversity that can be further improved with the proposed adaptive approach.
D.6 Controllable generation
For both tasks, we use the adaptive refinement strategy and set the rollback value $\sigma$ to $0.5$. We perform $5$ steps for the student generation and $10$ steps for the refinement with the UniPC solver. The cut-off thresholds correspond to $70$ and $50$ ImageReward percentiles for the mask-guided and edge-guided generation, respectively. We select random $600$ image-text pairs from the COCO2014 validation set for the edge-guided generation. For the mask-guided generation, we use $600$ semantic segmentation masks from the ADE20K dataset [59] and use the category names as the text prompts. For evaluation, we conduct the human study similar to D.1.
D.7 ImageReward inference costs
[…]
D.8 Additional visualizations
In Figure 23, 24 we present qualitative verification of the first observation (the student sometimes outperforms its teacher according to the human evaluation). Figure 25 supports the second observation (the student wins are more likely where its samples differ from the teacher ones). In Figure 26 we demonstrate two adaptive strategies (refining and regeneration), the examples confirm that the refinement strategy improves the image fidelity and does not significantly alter the textual alignment, while the regeneration strategy may improve textual alignment.
Figures 27, 28, 29 provide more qualitative comparisons of our approach for different tasks.
Figure 19: Visual examples of similar (Left) and dissimilar (Right) teacher and student samples for SD1.5 (a) and Dreamshaper v7 (b).
Figure 20: Visual examples of similar (Left) and dissimilar (Right) teacher and student samples for the architecture-based distillation.
Figure 21: Analysis results for the architecture-based distillation.
Figure 22: Analysis results for the consistency distillation on Dreamshaper v7.
Figure 23: Additional examples where the student (CD-SD1.5) outperforms its teacher (SD1.5) according to the human evaluation.
Figure 24: Additional examples where the student (CD-SDXL) outperforms its teacher (SDXL) according to the human evaluation.
Figure 25: Visualization of the student (CD-SDXL) and teacher (SDXL) samples for low and high distance ranges. The green outline corresponds to wins, while the red one - losses.
Figure 26: Visual examples of the refining and regeneration adaptive strategies.
Figure 27: Additional image editing results produced with our approach.
Figure 28: Additional results on Canny edge guided image generation with our approach.
Figure 29: Additional results on segmentation mask guided image generation with our approach.
Figure 30: Visual examples generated with various distilled text-to-image models for different seed values. CD-based students generate more diverse images than ADD-XL.